OpenStudy (anonymous):

Use a transformation to solve the equation. z – (–0.3) = –2.4 A. z = –2.7 B. z = –2.1 C. z = 2.1 D. z = 2.7

Mehek (mehek14):

– (–0.3) = +0.3 z + 0.3 = -2.4 -2.4 - 0.3 = ?

OpenStudy (anonymous):

thx ur awesome